If log a ≈ 0.250 and log b ≈ 0.644, then log b/a ≈
a. 2.576
b. 0.161
c. 0.394
d. There is insufficient information for a meaningful answer

Answer by jim_thompson5910(35256)

log (b/a)=log(b)-log(a)=0.644-0.250=0.394


So log (b/a)≈0.394